I have the answer to the equation I just need to figure out what the equation is. We are working on Solving Open sentences involving absolute value.
I just need the problem,the answer is -1 and 3.
thank you
Lori Taylor

Answer by scott8148(6628)

abs(x + a) = b

3 + a = b

-1 + a = -b

substituting ___ -1 + a = -3 - a ___ 2a = -2 ___ a = -1

substituting ___ 3 + (-1) = b ___ 2 = b

abs(x - 1) = 2